# On the Ability of Deep Networks to Learn Symmetries from Data: A Neural Kernel Theory

### v1

@@ B.3 Architecture trained on multiple seeds and multiple classes of rotated-MNIST
Appendix C Extension of the spectral theory to equivariant architectures
@@
$g_{T}.x\left(\begin{bmatrix}i_{x}\\ i_{y}\end{bmatrix}\right)=x\left(\begin{bmatrix}(i_{x}+1)\bmod n\\ i_{y}\end{bmatrix}\right).$ (63)
@@
$g_{R}.x\left(\begin{bmatrix}i_{x}\\ i_{y}\end{bmatrix}\right)=x\left(\begin{bmatrix}i_{y}\\ n-i_{x}\end{bmatrix}\right).$ (65)
@@
$f_{\text{FC}}(x)=A\frac{1}{\sqrt{k}}\phi(B\circledast x)_{v},$ (66)
@@
Let $K\in\mathbb{R}^{n^{2}\times n^{2}}$ denote the Conv-NNGP for a 1-layer fully convolutional network operating on a pair of images $x,x^{\prime}$ of size $n\times n$. This kernel contains an entry for every coordinate quadruplet $(i_{x},i_{y},i_{x}^{\prime},i_{y}^{\prime})$ across the images $x,x^{\prime}$. On the other hand, the Conv-NNGP of the network above $f_{\text{FC}}$ reduces this representation to a kernel of size $K_{\text{FC}}\in\mathbb{R}$, by letting $K_{\text{FC}}=\operatorname{Tr}(K)$ (Arora et al.,, 2019):
$K_{\text{FC}}(x,x^{\prime})$ $=\sum_{i_{x},i_{y}}K_{i_{x},i_{y},i_{x},i_{y}}(x,x^{\prime})$ (67)
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x},i_{y}}\cdot x^{\prime})\Big{\rangle}_{\Theta}$ (68)
@@
$f_{\text{GAP}}(x)=\frac{1}{\sqrt{k}n^{2}}\sum_{k}\sum_{i_{x}}\sum_{i_{y}}A_{1k}\phi(B_{k,i_{x},i_{y}}\cdot x)$ (69)
@@
$K_{\text{GAP}}(x,x^{\prime})$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}K_{i_{x},i_{y},i_{x}^{\prime},i_{y}^{\prime}}(x,x^{\prime})$ (70)
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot x^{\prime})\Big{\rangle}_{\Theta}$ (71)
Proposition 1.
@@
$K_{\text{FC}}(g_{T}.x,g_{T}.x^{\prime})$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot g_{T}.x),\phi(B_{k,i_{x},i_{y}}\cdot g_{T}.x)\Big{\rangle}_{\Theta}$ (72)
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(g_{T}^{-1}.B_{k,i_{x},i_{y}}\cdot x),\phi(g_{T}^{-1}.B_{k,i_{x},i_{y}}\cdot x)\Big{\rangle}_{\Theta}$ (73)
@@
$K_{\text{FC}}(g_{T}\cdot x,g_{T}\cdot x^{\prime})$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x}-1\bmod n}\cdot x),\phi(B_{k,i_{x}-1\bmod n,i_{y}}\cdot x)\Big{\rangle}_{\Theta}$ (74)
@@
$K_{\text{FC}}(g_{T}.x,g_{T}.x^{\prime})$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x},i_{y}}\cdot x)\Big{\rangle}_{\Theta}$ (75)
$=K_{\text{FC}}(x,x^{\prime})$ (76)
@@
$K_{\text{FC}}(x,g_{T}.x)$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x},i_{y}}\cdot g_{T}.x)\Big{\rangle}_{\Theta}$ (77)
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(g_{T}^{-1}.B_{k,i_{x},i_{y}}\cdot x)\Big{\rangle}_{\Theta}$ (78)
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x}-1\bmod n,i_{y}}\cdot x)\Big{\rangle}_{\Theta}$ (79)
@@
$K_{\text{GAP}}(x,g_{T}.x)$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot(g_{T}.x))\Big{\rangle}_{\Theta}$ (81)
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(g_{T}^{-1}.B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot(x))\Big{\rangle}_{\Theta}$ (82)
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x}^{\prime}-1\bmod n,i_{y}^{\prime}}\cdot(x))\Big{\rangle}_{\Theta}$ (83)
@@
$K_{\text{GAP}}(x,g_{T}.x)$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot x)\Big{\rangle}_{\Theta}$ (84)
$=K_{\text{GAP}}(x,x)$ (85)
@@
Proof.
Now we show that the kernel of a fully connected convolutional network is circulant over an orbit of image rotations. Consider a pair of images from this orbit $(x,x^{\prime})\in O_{R}$. We show next that applying the same image rotation $g_{R}$ to both these images leaves the kernel unaffected:
$K_{\text{FC}}(g_{R}.x,g_{R}.x^{\prime})$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot g_{R}.x),\phi(B_{k,i_{x},i_{y}}\cdot g_{R}.x)\Big{\rangle}_{\Theta}$ (86)
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(g_{R}^{-1}.B_{k,i_{x},i_{y}}\cdot x),\phi(g_{R}^{-1}.B_{k,i_{x},i_{y}}\cdot x)\Big{\rangle}_{\Theta}$ (87)
@@
$K_{\text{FC}}(g_{R}\cdot x,g_{R}\cdot x^{\prime})$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B^{\prime}_{k,n-i_{y},i_{x}}\cdot x),\phi(B^{\prime}_{k,n-i_{y},i_{x}}\cdot x)\Big{\rangle}_{\Theta}$ (88)
@@
$K_{\text{FC}}(g_{R}.x,g_{R}.x^{\prime})$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}^{\prime}\cdot x),\phi(B_{k,i_{x},i_{y}}^{\prime}\cdot x)\Big{\rangle}_{\Theta}$ (89)
@@
$K_{\text{FC}}(g_{R}.x,g_{R}.x^{\prime})$ $=K_{\text{FC}}(x,x^{\prime})$ (90)
@@
Another question of interest is, is there anything special about the kernel of two images belonging to the same orbit. Consider an image $x$ and its transformation $g_{R}.x$.
$K_{\text{FC}}(x,g_{R}.x)$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x},i_{y}}\cdot g_{R}.x)\Big{\rangle}_{\Theta}$ (91)
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(g_{R}^{-1}.B_{k,i_{x},i_{y}}\cdot x)\Big{\rangle}_{\Theta}$ (92)
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B^{\prime}_{k,n-i_{y},i_{x}}\cdot x)\Big{\rangle}_{\Theta}$ (93)
@@
A similar argument as in the case of shifts can be made to prove that the kernel is not, in general, rank-deficient. This can be obtained by choosing a “Dirac-like” image whose only active pixel is not at the center of the image. ∎
Proposition 4.
@@
Proof.
We study how changing the last layer from a fully connected to a global average pooling layer affects the results of the previous proposition.
$K_{\text{GAP}}(g_{R}.x,g_{R}.x^{\prime})$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot g_{R}.x),\phi(B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot g_{R}.x^{\prime})\Big{\rangle}_{\Theta}$ (95)
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(g_{R}^{-1}.B_{k,i_{x},i_{y}}\cdot x),\phi(g_{R}^{-1}.B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot x^{\prime})\Big{\rangle}_{\Theta}$ (96)
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B^{\prime}_{k,n-i_{y},i_{x}}\cdot x),\phi(B^{\prime}_{k,n-i_{y}^{\prime},i_{x}^{\prime}}\cdot x^{\prime})\Big{\rangle}_{\Theta}$ (97)
@@
$K_{\text{GAP}}(g_{R}.x,g_{R}.x^{\prime})$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B^{\prime}_{k,i_{x},i_{y}}\cdot x),\phi(B^{\prime}_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot x^{\prime})\Big{\rangle}_{\Theta}$ (98)
$=K_{\text{GAP}}(x,x^{\prime})$ (99)
@@
$K_{\text{GAP}}(x,g_{R}.x)$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot g_{R}.x)\Big{\rangle}_{\Theta}$ (100)
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(g_{R}^{-1}.B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot x)\Big{\rangle}_{\Theta}$ (101)
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B^{\prime}_{k,n-i_{y}^{\prime},i_{x}^{\prime}}\cdot x)\Big{\rangle}_{\Theta}$ (102)
@@
$K_{\text{GAP}}(x,g_{R}.x)$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B^{\prime}_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot x)\Big{\rangle}_{\Theta}$ (103)
$\neq K_{\text{GAP}}(x,x)$ (104)
In general, the kernel is not be constant over an orbit.
To prove that this kernel over an orbit is in general not rank deficient, one can resort to using a seed image that is the sum of two “Dirac-like” images, each with a single nonzero pixel, and both of them in an area contained within the filter size. ∎
Appendix D Varying the number of angles in an orbit

### v2

@@ D.3 Architecture Trained on Multiple Seeds and Multiple Classes of Rotated-MNIST
Appendix E Extension of the Spectral Theory to Equivariant Architectures
@@
$g_{T}.x\left(\begin{bmatrix}i_{x}\\ i_{y}\end{bmatrix}\right)=x\left(\begin{bmatrix}(i_{x}+1)\bmod n\\ i_{y}\end{bmatrix}\right).$
@@
$g_{R}.x\left(\begin{bmatrix}i_{x}\\ i_{y}\end{bmatrix}\right)=x\left(\begin{bmatrix}i_{y}\\ n-i_{x}\end{bmatrix}\right).$
@@
$f_{\text{FC}}(x)=A\frac{1}{\sqrt{k}}\phi(B\circledast x)_{v},$
@@
Let $K\in\mathbb{R}^{n^{2}\times n^{2}}$ denote the Conv-NNGP for a 1-layer fully convolutional network operating on a pair of images $x,x^{\prime}$ of size $n\times n$. This kernel contains an entry for every coordinate quadruplet $(i_{x},i_{y},i_{x}^{\prime},i_{y}^{\prime})$ across the images $x,x^{\prime}$. On the other hand, the Conv-NNGP of the network above $f_{\text{FC}}$ reduces this representation to a kernel of size $K_{\text{FC}}\in\mathbb{R}$, by letting $K_{\text{FC}}=\operatorname{Tr}(K)$ (Arora et al., 2019):
$K_{\text{FC}}(x,x^{\prime})$ $=\sum_{i_{x},i_{y}}K_{i_{x},i_{y},i_{x},i_{y}}(x,x^{\prime})$
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x},i_{y}}\cdot x^{\prime})\Big{\rangle}_{\Theta}$
@@
$f_{\text{GAP}}(x)=\frac{1}{\sqrt{k}n^{2}}\sum_{k}\sum_{i_{x}}\sum_{i_{y}}A_{1k}\phi(B_{k,i_{x},i_{y}}\cdot x)$
@@
$K_{\text{GAP}}(x,x^{\prime})$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}K_{i_{x},i_{y},i_{x}^{\prime},i_{y}^{\prime}}(x,x^{\prime})$
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot x^{\prime})\Big{\rangle}_{\Theta}$
Proposition 11
@@
$K_{\text{FC}}(g_{T}.x,g_{T}.x^{\prime})$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot g_{T}.x),\phi(B_{k,i_{x},i_{y}}\cdot g_{T}.x)\Big{\rangle}_{\Theta}$
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(g_{T}^{-1}.B_{k,i_{x},i_{y}}\cdot x),\phi(g_{T}^{-1}.B_{k,i_{x},i_{y}}\cdot x)\Big{\rangle}_{\Theta}$
@@
$K_{\text{FC}}(g_{T}\cdot x,g_{T}\cdot x^{\prime})$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x}-1\bmod n}\cdot x),\phi(B_{k,i_{x}-1\bmod n,i_{y}}\cdot x)\Big{\rangle}_{\Theta}$
@@
$K_{\text{FC}}(g_{T}.x,g_{T}.x^{\prime})$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x},i_{y}}\cdot x)\Big{\rangle}_{\Theta}$
$=K_{\text{FC}}(x,x^{\prime})$
@@
$K_{\text{FC}}(x,g_{T}.x)$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x},i_{y}}\cdot g_{T}.x)\Big{\rangle}_{\Theta}$
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(g_{T}^{-1}.B_{k,i_{x},i_{y}}\cdot x)\Big{\rangle}_{\Theta}$
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x}-1\bmod n,i_{y}}\cdot x)\Big{\rangle}_{\Theta}$
@@
$K_{\text{GAP}}(x,g_{T}.x)$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot(g_{T}.x))\Big{\rangle}_{\Theta}$
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(g_{T}^{-1}.B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot(x))\Big{\rangle}_{\Theta}$
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x}^{\prime}-1\bmod n,i_{y}^{\prime}}\cdot(x))\Big{\rangle}_{\Theta}$
@@
$K_{\text{GAP}}(x,g_{T}.x)$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot x)\Big{\rangle}_{\Theta}$
$=K_{\text{GAP}}(x,x)$
@@
Proof  Now we show that the kernel of a fully connected convolutional network is circulant over an orbit of image rotations. Consider a pair of images from this orbit $(x,x^{\prime})\in O_{R}$. We show next that applying the same image rotation $g_{R}$ to both these images leaves the kernel unaffected:
$K_{\text{FC}}(g_{R}.x,g_{R}.x^{\prime})$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot g_{R}.x),\phi(B_{k,i_{x},i_{y}}\cdot g_{R}.x)\Big{\rangle}_{\Theta}$
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(g_{R}^{-1}.B_{k,i_{x},i_{y}}\cdot x),\phi(g_{R}^{-1}.B_{k,i_{x},i_{y}}\cdot x)\Big{\rangle}_{\Theta}$
@@
$K_{\text{FC}}(g_{R}\cdot x,g_{R}\cdot x^{\prime})$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B^{\prime}_{k,n-i_{y},i_{x}}\cdot x),\phi(B^{\prime}_{k,n-i_{y},i_{x}}\cdot x)\Big{\rangle}_{\Theta}$
@@
$K_{\text{FC}}(g_{R}.x,g_{R}.x^{\prime})$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}^{\prime}\cdot x),\phi(B_{k,i_{x},i_{y}}^{\prime}\cdot x)\Big{\rangle}_{\Theta}$
@@
$K_{\text{FC}}(g_{R}.x,g_{R}.x^{\prime})$ $=K_{\text{FC}}(x,x^{\prime})$
@@
Another question of interest is whether the kernel is constant or rank-deficient over pairs of images belonging to the same orbit. Consider an image $x$ and its transformation $g_{R}.x$.
$K_{\text{FC}}(x,g_{R}.x)$ $=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x},i_{y}}\cdot g_{R}.x)\Big{\rangle}_{\Theta}$
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(g_{R}^{-1}.B_{k,i_{x},i_{y}}\cdot x)\Big{\rangle}_{\Theta}$
$=\sum_{i_{x},i_{y}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B^{\prime}_{k,n-i_{y},i_{x}}\cdot x)\Big{\rangle}_{\Theta}$
@@
A similar argument as in the case of shifts can be made to prove that the kernel is not, in general, rank-deficient. This can be proven by choosing a “Dirac-like” image whose only active pixel is not at the center of the image.
Proposition 14
@@
Proof  We study how changing the last layer from a fully connected to a global average pooling layer affects the results of the previous proposition.
$K_{\text{GAP}}(g_{R}.x,g_{R}.x^{\prime})$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot g_{R}.x),\phi(B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot g_{R}.x^{\prime})\Big{\rangle}_{\Theta}$
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(g_{R}^{-1}.B_{k,i_{x},i_{y}}\cdot x),\phi(g_{R}^{-1}.B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot x^{\prime})\Big{\rangle}_{\Theta}$
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B^{\prime}_{k,n-i_{y},i_{x}}\cdot x),\phi(B^{\prime}_{k,n-i_{y}^{\prime},i_{x}^{\prime}}\cdot x^{\prime})\Big{\rangle}_{\Theta}$
@@
$K_{\text{GAP}}(g_{R}.x,g_{R}.x^{\prime})$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B^{\prime}_{k,i_{x},i_{y}}\cdot x),\phi(B^{\prime}_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot x^{\prime})\Big{\rangle}_{\Theta}$
$=K_{\text{GAP}}(x,x^{\prime})$
@@
$K_{\text{GAP}}(x,g_{R}.x)$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot g_{R}.x)\Big{\rangle}_{\Theta}$
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(g_{R}^{-1}.B_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot x)\Big{\rangle}_{\Theta}$
$=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B^{\prime}_{k,n-i_{y}^{\prime},i_{x}^{\prime}}\cdot x)\Big{\rangle}_{\Theta}$
@@
$K_{\text{GAP}}(x,g_{R}.x)$ $=\sum_{\begin{subarray}{c}i_{x},i_{y}\\ i_{x}^{\prime},i_{y}^{\prime}\end{subarray}}\sum_{k}\Big{\langle}\phi(B_{k,i_{x},i_{y}}\cdot x),\phi(B^{\prime}_{k,i_{x}^{\prime},i_{y}^{\prime}}\cdot x)\Big{\rangle}_{\Theta}$
$\neq K_{\text{GAP}}(x,x)$
In general, the kernel is not constant over an orbit.
To prove that this kernel over an orbit is in general not rank deficient, one can resort to using a seed image that is the sum of two “Dirac-like” images, each with a single nonzero pixel, and both of them in an area contained within the filter size.
Appendix F Varying the Number of Angles in an Orbit